OpenStudy (anonymous):

PLEASE HELP WITH THESE TWO QUESTIONS 1...What are the values of a, b, and c in the function? List them in order with commas, such as 1,-2,3. y = -2x2 + 5x - 4 2...What is the vertex for the function? (You will want to find the axis of symmetry first.) Write as an ordered pair, leaving no spaces, such as (1,-4). y = -3x2 + 6x

OpenStudy (anonymous):

1. -2, 5, -4

OpenStudy (anonymous):

Do you know the second one?

OpenStudy (anonymous):

\[Xv= (-b)/2a \] \[Yv=-\Delta/4a\]

OpenStudy (anonymous):

x= -b/(2a) y(-b/2a)=

OpenStudy (anonymous):

\[\Delta=b^2-4ac\]

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

no i dont know the anwser ? lol

OpenStudy (anonymous):

just put the numbers into the formulas and you'll get the answers..

OpenStudy (anonymous):

the answer is ( 1,3)

OpenStudy (anonymous):

\[a=-3 , b=6, x=(-6)/2(-3)=1, y(1)= -3(1)^2+6(1)=3\]

OpenStudy (anonymous):

So the vertex is (1,3)